tod so 48 An IV infusion order is received for lidocaine (Xvlocaine) 2g in 250mL of D5W, to be infused at 3mg/min. What will 48 and 49
Answer #1

48. Lidocaine 2 gms in 250 ml = 2000 mg in 250 ml

The amount of solution required is = Required dose \times available volume \div available dose

                                                       = 3 mg \times 250 ml \div 2000 mg            

                                                       = 0.375 ml per minute

The flow rate per hour = 0.375 \times 60 = 22.5 ml/ hour

Option b is the correct answer.

49. KCL 20 mEq in 1000 ml of D5W.

In 1 ml of solution, the dose is = 20 \div 1000 = 0.02 mEq per ml

In 1 hour the KCL received = 125 \times 0.02 = 2.5 mEq per hour

Option b is the correct answer.